** Estimate the energy splitting between the lowest two levels for an electron in a three-dimensional cube-shaped potential box with the side length 10 nm.

1◦ 1 meV

2◦ 10 meV

3◦ 100 meV

4◦ 0.11 eV

5◦ 0.22 eV

6◦ 1.12 eV


1
Expert's answer
2021-07-19T09:48:31-0400

Energy in ground state, "E_1=\dfrac{(1)^2\pi^2\hbar^2}{2m_eL^2}"

"E_1=\dfrac{\pi^2(1.054\times10^{-34})^2}{2(9.1\times10^{-31})(10\times10^{-9})^2}"

"E_1=6.022\times10^{-22}\space J"


Energy in first excited state, "E_2=\dfrac{(2)^2\pi^2\hbar^2}{2m_eL^2}=4\times6.022\times10^{-22}\space J"

Energy difference, "\Delta E=E_2-E_1=3\times6.022\times10^{-22}\space J"

"\Delta E=112.82\times10^{-3}\space eV=0.11\space eV"